OpenStudy (anonymous):

check my answer..... The area of a rectangular athletic field is represented by the expression 49y^2-9z^2 square meters. Write an algebraic expression to represent one possible set of dimensions (in the sense “length times width”) of the athletic field. Include correct units with your solution.

OpenStudy (anonymous):

A = L x W sq. meters LxW = 49y^2 - 9z^2 sq. meters LxW = (7y +3z)(7y - 3z) sq. meters L = 7y + 3z meters W = 7y -3z meters

OpenStudy (anonymous):

perfection mang!